Vector B has x, y, and z components of 4.00, 6.00, and 3.00 units, respectively. Calculate the magnitude

of B and the angles that B makes with the coordinate axes.


1
Expert's answer
2021-10-01T12:22:19-0400
B=Bx2+By2+Bz2=42+62+82=10.8.θx=arccosBxB=68°, θy=arccosByB=56°, θz=arccosBzB=42°.|B|=\sqrt{B_x^2+B_y^2+B_z^2}=\sqrt{4^2+6^2+8^2}=10.8.\\ \theta_x=\arccos\frac{B_x}{|B|}=68°,\\\space\\ \theta_y=\arccos\frac{B_y}{|B|}=56°,\\\space\\ \theta_z=\arccos\frac{B_z}{|B|}=42°.